Westinghouse wants to determine the lifetime of its standard lightbulbs with an error 13 hours. Standard deviation is 120 hours. What sample size should they have to be 95% confident of the lifetime of their standard lightbulb?

a) 18.09 lightbulbs

b) 19 lightbulbs

c) 328 lightbulbs

d) 327.33 lightbulbs
Answer #1

Solution :

Given that,

standard deviation = = 120

margin of error = E = 13

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

Sample size = n = ((Z/2 * ) / E)2

= ((1.96 * 120) / 13)2

= 327.33

Sample size = 327.33 lightbulbs

option D) is correct